OpenStudy (anonymous):

(x-2)^(1/3)=5 how would I find x in this equation?

OpenStudy (anonymous):

You cube both sides to get rid of the cubic root. Then you add 2 to both sides.

OpenStudy (anonymous):

\[5 = (x - 2)^{\frac{1}{3}}\]\[5^3 = x - 2\]\[125 = x - 2\]\[127 = x\]